OpenStudy (anonymous):

A spherical water balloon has a radius of 3 inches. How many cubic inches of water will it hold?

OpenStudy (bradely):

Volume ,V = (4/3) πr^3 =(4/3) π(3)^3 =36π cubic inches source:www.mathskey.com